OpenStudy (anonymous):

determine the truth value of the statement 15 − 7 = 22 or 4 + 9 = 13, and 9 − 8 = 1. Substitute the truth values for p, q, and r to determine the truth value for the letter’s claim (p /\ q) --> r (F /\ T) --> F F --> F T

OpenStudy (jiteshmeghwal9):

15+7*

OpenStudy (anonymous):

this is the answer 15+7*

OpenStudy (anonymous):

I need the truth value

OpenStudy (amistre64):

place T or F in place of those equations for starters

OpenStudy (amistre64):

(__ v __) ^ __

OpenStudy (anonymous):

can you you me what to input for the answer

OpenStudy (amistre64):

im not going to do the work for you ; no.

OpenStudy (amistre64):

you seem to have edited it tho .... sneaky

OpenStudy (amistre64):

im not sure how the extra stuff fits in to the picture tho; it doesnt fit the question

OpenStudy (anonymous):

my instructor said I did it wrong the first time and gave me the addiional information to hlep me

OpenStudy (amistre64):

(15 − 7 = 22) or (4 + 9 = 13), and (9 − 8 = 1) lets start by defining the truth value of these statements 15 − 7 = 22 ; True or False? 4 + 9 = 13 ; True or False? 9 − 8 = 1 ; True or False?

OpenStudy (anonymous):

false,true,true. I got that part I jus need to input the other

OpenStudy (amistre64):

then lets replace them with their truth values (15 − 7 = 22) or (4 + 9 = 13), and (9 − 8 = 1) ( F or T ) and T and write it in notation (F v T) ^ T just like the order of operations, we can do whats inside the (...) first F v T ; does this gives us a True or False result?

OpenStudy (anonymous):

(F^T)->T would that be the way

OpenStudy (amistre64):

no

OpenStudy (amistre64):

the arrow replaces the word "then", it points to a consequence there are no arrows to be had ... yet

OpenStudy (anonymous):

So that would give us a true value. Right ?

OpenStudy (amistre64):

correct; FvT uses an "or" operator. or is true when at least one of them is true (FvT) = T lets replace that then (F v T) ^ T T ^ T so, can you tell me the truth value of whats left? T ^ T = true or false?

OpenStudy (anonymous):

false

OpenStudy (amistre64):

lets review the and operator (^) is only true with all of its arguments are true since we have 2 arguments, both are true, we are using the "and" operator ... this tells us that: T^T = T since we are left with only one arguement, there is nothing left to determine .....

OpenStudy (amistre64):

(15 − 7 = 22) or (4 + 9 = 13), and (9 − 8 = 1) ( F or T ) and T T and T T

OpenStudy (amistre64):

sometimes it really is better just to leave the words inplace :)